OpenStudy (anonymous):

A right triangle has legs with lengths 7 cm and 24 cm. What is the length of the hypotenuse? ____ cm?

OpenStudy (mrnood):

this is the very basis of understanding traingles and oyu should work this out for yourself The equaiotn you need is the Pythagoras eqution ans is the sum of th esqures of the 2 legs is equal to the square of the hypotenuse SO write down 7^2 =?? 24^2 =?? 7^2+24^2 = ?? That last value is then the sqare of the hypotenuse - so take the square root and you will hav eth eanswer: 

OpenStudy (anonymous):

I got 625 when I did \[7^{2} + 24^{2}\]

OpenStudy (mrnood):

that's good - and correct now take sqrt (625)

OpenStudy (anonymous):

I got 25. :)

OpenStudy (mrnood):

Perfectly correct just remember the hypotenuse is the longest side and a^2 + b^2 = c^2

OpenStudy (mrnood):